A syringe with a maximum volume of 60.0 cm cubed takes in a 14.5 cm cubed sample of oxygen gas at 23.3 degrees celsius. What maximum variation in temperature does the oxygen undergo before the piston is completely pushed out of the syringe
1
Expert's answer
2020-11-26T05:09:49-0500

From Charles law expression of an ideal gas;

"V_1\over T_1" "=" "V_2\over T_2"

"V_1=14.5cm^3"

"V_2=60.0cm^3"

"T_1=23\u00b0C"

"T_2=unknown"

"\therefore" "T_2=" "V_2\times T_1\over V_1"

"T_2=" "60.0cm^3\times 23.0°C\over 14.5cm^3"

"T_1=95.17\u00b0C"

The variation in temperature"=95.17\u00b0C-14.5\u00b0C=80.67\u00b0C"